Crosby Department Store estimates inventory by using the retail inventory method. The following information was developed:

At Cost At Retail
Beginning inventory $465,000 $ 750,000
Goods purchased 900,000 1,350,000
Net sales 1,500,000


The estimated cost of the ending inventory is

$210,000.
$372,000.
$400,000.
$390,000.
Answer #1

Goods available at cost = 465,000+900,000 = 1,365,000

At retail = 750,000+1,350,000 = 2,100,000

Cost = 1,365,000/2,100,000 = 65% of retail

Cost of Goods Sold = 1,500,000 * 65% = 975,000

Ending inventory = cost of goods available for sale - Cost of Goods Sold

= (465,000+900,000) - 975,000

= 390,000